# Tony Lindeberg

### v1

@@ Abstract
By comparison to results of neurophysiological measurements of direction and speed selectivity for biological neurons in the primary visual cortex, we find that our theoretical results are qualitatively consistent with (i) velocity-tuned visual neurons that are sensitive to particular motion directions and speeds, and (ii) different visual neurons having broader vs. sharper direction and speed selective properties.
@@ 1 Introduction
Regarding the analysis of visual motion, it is in particular essential to understand the properties of the spatio-temporal receptive fields in the early visual pathway. The functional properties of the receptive fields of the visual neurons can aptly be regarded as the main primitives in computational modelling of the visual system. Reconstructing a full spatio-temporal receptive field does, however, constitute a non-trivial task, since it constitutes an inverse problem, which requires measurements of the response properties of each visual neuron to a large number of stimuli, and also a model of the receptive field to relate those measurements to. Probing the direction and speed selectivity properties of a visual neuron is, on the other hand, a much more straightforward task, which can be accomplished by moving a probing stimulus with different motion directions and different motion speeds within the support region of the receptive field. Consequently, there are substantially more available results concerning direction selectivity properties of visual neurons in the primary visual cortex (V1) and the middle temporal visual area (MT) (Hubel 1959, Hubel and Wiesel 1962, Emerson and Gerstein 1977, Albright 1984, Albright et al. 1984, Ganz and Felder 1984, Mikami et al. 1986a; 1986b, Orban et al. 1986, Lagae et al. 1993, Orban 1997, Livingstone 1998, Livingstone and Conway 2003, Born and Bradley 2005, Churchland et al. 2005, Gur et al. 2005, Moore et al. 2005, Priebe and Ferster 2005, Priebe et al. 2010, Wang and Yao 2011, Dai et al. 2025, Khamiss et al. 2025) than there are publicly available reconstructed spatio-temporal receptive fields (DeAngelis et al. 1995; 2004, de Valois et al. 2000). For this reason, it is of interest to establish theoretically based relationships between the properties of the spatio-temporal receptive fields of visual neurons and their direction selectivity properties, in order to bridge the gap between these two ways of characterizing motion-sensitive neurons.
@@
The subject of this paper is to address this topic regarding the direction and speed selectivity properties of idealized mathematical models of spatio-temporal receptive fields in the primary visual cortex. For a theoretically well-founded way of modelling the spatio-temporal receptive fields of simple and complex cells for higher mammals, based on the generalized Gaussian derivative model for visual receptive fields (Lindeberg 2013; 2021), we will present an in-depth theoretical analysis of the direction and speed selectivity properties for idealized models of the spatio-temporal receptive fields of simple cells and complex cells based on this theory.
@@
A main motivation for using this model for visual receptive fields, is that it is theoretically well motivated in the sense that the shapes of the corresponding spatio-temporal receptive fields can be derived by necessity, from structural properties of the environment in combination with internal consistency requirements to guarantee consistent treatment of image structures over multiple spatial and temporal scales.
@@
The methodology that we will follow in this treatment is to subject idealized models of simple and complex cells according to the generalized Gaussian derivative theory to similar probing mechanisms, as are used for probing the direction and speed selectivity properties of biological neurons. Specifically, we will focus on variabilities in the shapes of the spatio-temporal receptive fields, as induced by spanning the degrees of freedom of geometric image transformations, and analyze how these variabilities in the shapes of the spatio-temporal receptive fields affect the direction selectivity properties. In this way, we will provide an framework for how observed variabilities in the direction and speed selectivity properties of biological neurons could be explained in terms of variabilities in the shapes of the underlying simple cells, as well as the use of such simple cell primitives in our idealized models of complex cells.
@@ 5.1 Analysis for integrated velocity-tuned models of complex cells
By summarizing results for the direction selectivity of the different idealized models of simple cells in Section 4, we find that provided that the spatial and the temporal scale parameters $\sigma_{1}$ and $\sigma_{t}$ are coupled to the velocity parameter $v$ of the spatio-temporal receptive field according to $\sigma_{1}/\sigma_{t}=v$, the combined direction and speed selectivity is of the form
@@ 7.1 Relations to neurophysiological measurements of motion-sensitive neurons in the primary visual cortex
Given such an expansion of the shapes of the shapes of the spatio-temporal receptive fields in the primary visual cortex over the degrees of freedom of local Galilean transformations, a consequence of that would then be variabilities in the direction selectivity properties and the speed sensitivity properties of the visual neurons, as recorded in neurophysiological measurements.
@@ 8 Suggestions for further neurophysiological experiments
Answering these questions would provide further cues towards understanding the computational function of the spatio-temporal receptive fields in the primary visual cortex, and concerning to what extent we can regard the influence of geometric image transformations on the receptive field responses as a primary factor for the development of the visual receptive fields. Such more detailed comparisons between the properties of the theoretical model and the properties of the biological neurons could also provide cues to possible ways of extending the theoretical model for spatio-temporal receptive fields with additional [^4] The current model for linear spatio-temporal receptive fields, that we base this work upon, has been derived under minimal assumptions regarding the structure of the environment in combination with internal consistency requirements to guarantee theoretically well-founded treatment of image structures over multiple spatial and temporal scales (Lindeberg 2011; 2021). a priori information, if motivated from results from neurophysiological experiments.
@@ 9 Summary and discussion
By subjecting such idealized models of motion-sensitive neurons in the primary visual cortex to a structurally similar probing method as used for probing the direction and speed selectivity properties of visual neurons, we have derived closed form expressions for the direction and speed selectivity properties of our idealized models of simple cells, as summarized in Equations (50) and (51) for the special case when spatial and temporal scale parameters $\sigma_{1}$ and $\sigma_{t}$ are coupled to the speed parameter $v$ of the spatio-temporal receptive field according to $\sigma_{1}/\sigma_{t}=v$. We have also derived more general results for such combined direction and speed selectivity properties, reproduced in Figure 10 and Equation (52) for the more general case, when there is no coupling between the scale parameters in relation to the speed parameter of the receptive field.
@@
By comparisons with the results of neurophysiological recordings of motion-sensitive neurons in the primary visual cortex in Section 7.1, the presented theoretical results are qualitatively consistent with the previously reported class of velocity-tuned motion sensitive neurons, in the sense that the models of the neurons respond maximally to the combination of a preferred motion direction and a preferred motion speed, and that the magnitude of the response decreases when changing the motion direction and/or the motion speed from the preferred value.
By comparisons with overall results concerning direction selective neurons in the primary visual cortex V1 and the middle temporal area MT in Section 7.2, with their organization into subcolumns preferring motion in different directions, we have also found this organization consistent with the hypothesis proposed in Lindeberg (2023b; 2025d), that the population of receptive fields of the simple cells in the primary visual cortex ought to be covariant under local Galilean transformations.
This would then specifically imply that the spatio-temporal receptive ought to have their shapes expanded over the degrees of freedom of local transformations. In practice, this would imply that there would be multiple copies of similar types of spatio-temporal receptive fields for different values of the velocity parameter $v$ in the idealized model (3.1) of simple cells, as illustrated in Figures 2 and 2.

### v2

@@ Abstract
By comparison to results of neurophysiological measurements of direction and speed selectivity for biological neurons in the primary visual cortex, we find that our theoretical results are consistent with the properties that (i) velocity-tuned visual neurons are sensitive to particular motion directions and speeds, and (ii) different visual neurons having broader vs. sharper direction and speed selective properties.
@@ 1 Introduction
Regarding the analysis of visual motion, it is in particular essential to understand the properties of the spatio-temporal receptive fields in the early visual pathway. The functional properties of the receptive fields of the visual neurons can aptly be regarded as the main primitives in computational modelling of the visual system. Reconstructing a full spatio-temporal receptive field does, however, constitute a non-trivial task, since it constitutes an inverse problem, which requires measurements of the response properties of each visual neuron to a large number of stimuli, and also a model of the receptive field to relate those measurements to. Probing the direction and speed selectivity properties of a visual neuron is, on the other hand, a much more straightforward task, which can be accomplished by moving a probing stimulus with different motion directions and different motion speeds within the support region of the receptive field. Consequently, there are substantially more available results concerning direction selectivity properties of visual neurons in the primary visual cortex (V1) and the middle temporal visual area (MT) (Hubel 1959, Hubel and Wiesel 1962, Emerson and Gerstein 1977, Albright 1984, Albright et al. 1984, Ganz and Felder 1984, Mikami et al. 1986a; 1986b, Orban et al. 1986, Lagae et al. 1993, Orban 1997, Livingstone 1998, Livingstone and Conway 2003, Born and Bradley 2005, Churchland et al. 2005, Gur et al. 2005, Moore et al. 2005, Priebe and Ferster 2005, Priebe et al. 2010, Wang and Yao 2011, Dai et al. 2025, Khamiss et al. 2025) than there are publicly available reconstructed spatio-temporal receptive fields (DeAngelis et al. 1995; 2004, de Valois et al. 2000). For this reason, it is of interest to establish theoretically based relationships between the properties of the spatio-temporal receptive fields of visual neurons and their direction selectivity properties, in order to bridge the explanatory gap between these two ways of characterizing the properties of motion-sensitive neurons.
The subject of this paper is to address this topic regarding the direction and speed selectivity properties of idealized mathematical models of spatio-temporal receptive fields in the primary visual cortex, which mimic the qualitative properties of biological motion-sensitive neurons, as have been experimentally recorded by Albright et al. (1984), Orban et al. (1986), Orban (1997), Movshon and Newsome (1996) and Churchland et al. (2005), see Born and Bradley (2005) and Elstrott and Feller (2009) for more extensive reviews and the forthcoming Section 7 for more specific relations between the theoretical results to be derived in this paper and previously established neurophysiological results regarding motion-sensitive neurons in areas V1 and MT
@@
For a theoretically well-founded way of modelling the spatio-temporal receptive fields of simple and complex cells for higher mammals, based on the generalized Gaussian derivative model for visual receptive fields (Lindeberg 2013; 2021), we will present an in-depth theoretical analysis of the direction and speed selectivity properties for idealized models of the spatio-temporal receptive fields of simple cells and complex cells based on this theory.
@@
A main motivation for using this model for visual receptive fields is that it is theoretically well motivated, in the sense that the shapes of the corresponding spatio-temporal receptive fields can be derived by necessity, from structural properties of the environment in combination with internal consistency requirements to guarantee consistent treatment of image structures over multiple spatial and temporal scales.
@@
The methodology that we will follow in this treatment is to subject idealized models of simple and complex cells according to the generalized Gaussian derivative theory to similar probing mechanisms, as are used for probing the direction and speed selectivity properties of biological neurons in neurophysiological experiments. Specifically, we will focus on variabilities in the shapes of the spatio-temporal receptive fields, as induced by spanning the degrees of freedom of geometric image transformations, and analyze how these variabilities in the shapes of the spatio-temporal receptive fields affect the direction selectivity properties. In this way, we will provide an framework for how observed variabilities in the direction and speed selectivity properties of biological neurons could be explained in terms of variabilities in the shapes of the underlying simple cells, as well as the use of such simple cell primitives in complementary idealized models of complex cells.
@@
An underlying motivation for the proposed new type of mathematical analysis of properties of computational models of motion-sensitive visual neurons is that a deeper understanding of their properties, including how these properties relate to neurophysiologically measurable entities of biological neurons, could allow for better computational models of motion-sensitive neurons, and thereby a better understanding of the functional properties of the early vision system.
@@ 5.1 Analysis for integrated velocity-tuned models of complex cells
By summarizing the results for the direction selectivity properties of the different idealized models of simple cells in Section 4, we find that provided that the spatial and the temporal scale parameters $\sigma_{1}$ and $\sigma_{t}$ are coupled to the velocity parameter $v$ of the spatio-temporal receptive field according to $\sigma_{1}/\sigma_{t}=v$, the combined direction and speed selectivity is of the form
@@ 7.1 Relations to neurophysiological measurements of motion-sensitive neurons in the primary visual cortex
Given such an expansion of the shapes of the spatio-temporal receptive fields in the primary visual cortex over the degrees of freedom of local Galilean transformations, a consequence of that would then be variabilities in the direction selectivity properties and the speed sensitivity properties of the visual neurons, as recorded in neurophysiological measurements.
@@ 8 Suggestions for further neurophysiological experiments
Answering these theoretically motivated questions would provide further cues towards understanding the computational function of the spatio-temporal receptive fields in the primary visual cortex, and concerning to what extent we can regard the influence of geometric image transformations on the receptive field responses as a primary factor for the development of the visual receptive fields. Such more detailed comparisons between the properties of the theoretical model and the properties of the biological neurons could also provide cues to possible ways of extending the theoretical model for spatio-temporal receptive fields with additional [^4] The current model for linear spatio-temporal receptive fields, that we base this work upon, has been derived under minimal assumptions regarding the structure of the environment in combination with internal consistency requirements to guarantee theoretically well-founded treatment of image structures over multiple spatial and temporal scales (Lindeberg 2011; 2021). a priori information, if motivated from results from additional neurophysiological experiments.
@@ 9 Summary and discussion
By subjecting such idealized models of motion-sensitive neurons in the primary visual cortex to a structurally similar probing method as used for probing the direction and speed selectivity properties of visual neurons in neurophysiological experiments, we have derived closed form expressions for the direction and speed selectivity properties of our idealized models of simple cells, as summarized in Equations (50) and (51) for the special case when spatial and temporal scale parameters $\sigma_{1}$ and $\sigma_{t}$ are coupled to the speed parameter $v$ of the spatio-temporal receptive field according to $\sigma_{1}/\sigma_{t}=v$. We have also derived more general results for such combined direction and speed selectivity properties, reproduced in Figure 10 and Equation (52) for the more general case, when there is no coupling between the scale parameters in relation to the speed parameter of the receptive field.
@@
By comparisons with the results of neurophysiological recordings of motion-sensitive neurons in the primary visual cortex, the presented theoretical results are consistent with the previously reported class of velocity-tuned motion sensitive neurons reported by Orban et al. (1986), in the sense that the models of the neurons respond maximally to the combination of a preferred motion direction and a preferred motion speed, and that the magnitude of the response decreases when changing the motion direction and/or the motion speed from the preferred value.
By comparisons with overall results concerning direction selective neurons in the primary visual cortex V1 and the middle temporal area MT in Section 7.2, as reported by Albright et al. (1984), see Born and Bradley (2005), Elstrott and Feller (2009) Orban et al. (1986), Orban (1997), Movshon and Newsome (1996) and Churchland et al. (2005), and as summarized in Section 7.1, with their organization into subcolumns preferring motion in different directions, we have also found this organization consistent with the hypothesis proposed in Lindeberg (2023b; 2025d), that the population of receptive fields of the simple cells in the primary visual cortex ought to be covariant under local Galilean transformations.
This would then specifically imply that the spatio-temporal receptive fields ought to have their shapes expanded over the degrees of freedom of local transformations. In practice, this would imply that there would be multiple copies of similar types of spatio-temporal receptive fields for different values of the velocity parameter $v$ in the idealized model (3.1) of simple cells, as illustrated in Figures 2 and 2.